OpenStudy (anonymous):

How do you get from -(2)/((1+x)^2)=-1 to (1+x)^2=2? Does multiplying both side by negative 2 satisfie this? Wouldn't the left side become (4)/((1+x)^2)? If you want the entire problem, I can post it, but basically, I had to derive a function, f, and got -(2)/((1+x)^2), and now I have to find the value, C guarenteed by the mean value theorm, so I set f'(x) = to the slope of the secant line. Now I'm trying to solve for it.

OpenStudy (anonymous):

How do I get \[-\frac{ 2 }{(1+x)^{2} }=-1\] to be \[(1+x)^{2}=2\] ? Multiply both sides by -2? Wouldn't that give me \[\frac{ 4 }{ (1+x)^{2} }=2\]

OpenStudy (anonymous):

The question is: Given the function \[f(x)=\frac{ 1-x }{ 1+x }\] what is the slope of the secant line through the points (0,f(0)) and (1,f(1))? Find the value c between 0 and 1 guaranteed by the mean value theorm. I got: f(0)=1 and f(1)=0, so Msec=-1. \[f'(x)=-\frac{ 2 }{ (1+x)^{2} }\] So \[-\frac{ 2 }{ (1+x)^{2} }=-1\] What to do from here?